OpenStudy (anonymous):

you select a card a random without replacing the card you select a second card find the probability P(M, then H) Mathemtics

OpenStudy (anonymous):

1/55

OpenStudy (kirbykirby):

what's M for?

OpenStudy (anonymous):

I'm trying to figure out how they got the answer I'm not really looking for answers to the problems. But someone to help me with the equation

OpenStudy (anonymous):

P(M, then H) = P(M) x P(H) =(2/7) x( 1/6 )= ? ^ ^ | only 1 card with letter H There are 2 cards with letter M

OpenStudy (kirbykirby):

what does the "M" stand for?

OpenStudy (anonymous):

Mathematics

OpenStudy (kirbykirby):

Mathematics card? What kind of deck of cards are you using?

OpenStudy (anonymous):

yes mathematics cards I'm Sorry

OpenStudy (kirbykirby):

I don;t understand what kind of cards these are lol sorry.

OpenStudy (anonymous):

its on one of my math tests

OpenStudy (anonymous):

and the full question is

OpenStudy (anonymous):

you select a card at random. without replacing the card you select a second card find the probability' MATHEMATICS

OpenStudy (kirbykirby):

oh the cards are the letters forming the word 'Mathematics' ?

OpenStudy (anonymous):

P(M,THEN H)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I don't know how they got 1/55

OpenStudy (kirbykirby):

Ok got it :)

OpenStudy (kirbykirby):

\[\frac{2}{11}*\frac{1}{10}\]

OpenStudy (kirbykirby):

You have 2 'M" out of 11 letters, so to get an M on the first turn, you have a probability of 2/11

OpenStudy (anonymous):

Ok I see now.

OpenStudy (anonymous):

so then you don't replace the card

OpenStudy (anonymous):

that makes H for the 1 out of 10. Right?

OpenStudy (kirbykirby):

yes exactly :)

OpenStudy (anonymous):

Thank You very Much

OpenStudy (kirbykirby):

yw